: The length of a rectangular garden is four times its width; if the area of the garden is 196 square meters, what are its dimensions?
Answer by checkley75(3666)

L=4W
LW=196
4W*W=196
4W^2=196
W^2=196/4
W=SQRT49
W=7 ANS FOR THE WIDTH.
L=4*7=28 ANS. FOR THE LENGTH.
pROOF:
7*28=196
196=196